# Breach in the Shield: Unveiling the Vulnerabilities of Large Language Models

### v1

@@
Abstract.
Large Language Models (LLMs) and Vision-Language Models (VLMs) have become essential to general artificial intelligence, exhibiting remarkable capabilities in task understanding and problem-solving. However, the real-world reliability of these models critically depends on their stability, which remains an underexplored area. Despite their widespread use, rigorous studies examining the stability of LLMs under various perturbations are still lacking. In this paper, we address this gap by proposing a novel stability measure for LLMs, inspired by statistical methods rooted in information geometry. Our measure possesses desirable invariance properties, making it well-suited for analyzing model sensitivity to both parameter and input perturbations. To assess the effectiveness of our approach, we conduct extensive experiments on models ranging in size from 1.5B to 13B parameters. Our results demonstrate the utility of our measure in identifying salient parameters and detecting vulnerable regions in input images or critical dimensions in token embeddings. Furthermore, leveraging our stability framework, we enhance model robustness during model merging, leading to improved performance.
Large Language Models, Stability Measure
1. Introduction
Figure 1. A case study of the Qwen-VL model on the SCI-QA. The image on the far right shows the per-pixel FI values calculated for this case, while the middle image illustrates how perturbing the top 10 positions with the highest FI values induces hallucination
Figure 2. A case study of the Qwen2 on the MMLU-geography. The figure illustrates the accuracy of responses to multiple-choice problems as an increasing percentage of parameters are zeroed out. “FI-High” denotes zeroing out parameters with high FI values, while “Random” represents randomly zeroing out parameters.
Large Language Models (LLMs) and Vision Language Models (VLMs) such as GPT (Brown et al., 2020) and Llama (Touvron et al., 2023), have revolutionized the field of Natural Language Processing (NLP), exhibiting remarkable proficiency across a variety of tasks (Achiam et al., 2023; Jiang et al., 2023) and modalities (Bai et al., 2023; Liu et al., 2024b). These modern LLMs are massive in size, trained on vast amounts of data, and meticulously aligned to prevent from generating harmful content (Perez et al., 2022), leaking private information (Zhang et al., 2024), or exhibiting sexual or religious bias (Xie and Lukasiewicz, 2023).
@@ 1. Introduction
Externally, LLMs are vulnerable to input perturbations, such as Embedding-Corrupted Prompts (Fort, 2023; Liu et al., 2024c). This susceptibility extends to visual inputs in VLMs, where adversarially optimized images can drastically alter model behavior (Qi et al., 2024). Beyond adversarial attacks, VLMs exhibit high sensitivity to perturbations in specific local regions of an image—a common issue, as user-uploaded images often suffer from blurring, masking, or low resolution.
Our case study on the Qwen-VL model (Bai et al., 2023) highlights this vulnerability. We applied our proposed stability measure to each pixel in the input image, as shown in the right columns of Figure 1. Based on these calculations, we masked the top 10 most sensitive patches irrelevant to the prompt question, which effectively misled the model into generating incorrect responses.
@@
Internally, LLM stability is further challenged by parameter perturbations, often introduced through quantization and model merging. While these techniques improve deployment efficiency by reducing inference costs (Frantar and Alistarh, 2023; Ashkboos et al., 2024), they can also induce hallucinations and degrade performance (Men et al., 2024; Yu et al., 2024; Li et al., 2024).
Not all parameters are equally susceptible. As shown in Figure 2, the model retains similar performance even when 5% of randomly selected parameters are dropped. However, zeroing out just 1% of the parameters with the highest Fisher Information (FI) can reduce accuracy below random guessing (0.25), highlighting the critical role of identifying these vulnerable locations.
@@
To address these challenges, we propose a novel stability measure called FI, First order local Influence, to quantitatively assess the stability of LLMs against local perturbations. Specifically, we construct a perturbation manifold that encompasses all perturbed models, along with its associated geometric properties. Our stability measure quantifies the degree of local influence of the perturbation to a given objective function within this manifold and thus reflects the stability of each component of LLMs.
@@
This measure is highly versatile, effectively assessing model stability under both external and internal perturbations across various levels of granularity—from individual parameters to input features such as pixels and patches.
Existing approaches, such as the Jacobian norm (Novak et al., 2018) and Saliency Map (Simonyan et al., 2013), can also serve as stability measures for large language models. We set these methods as competing baselines and conduct various numerical analyses to demonstrate the advantages of our approach. Moreover, among all these methods, ours is the only one that remains invariant to parameter reparameterization, a crucial property for interpretability, as discussed in Section 3.
@@
We conduct extensive studies to validate our method in identifying fragile pixels in vision inputs of VLMs (Subsections 4.1 and 4.2), vulnerable embedding dimensions of tokens in LLMs (Subsection 4.3), and salient parameters across models of varying sizes (Subsection 5.1).
Furthermore, we integrate our measure into various model merging methods (Subsection 5.2), demonstrating that protecting salient parameters with high FI values can significantly mitigate degradation caused by model merging, underscoring the critical role of FI in optimizing model performance.
Our main contributions are summarized as follows:
  1. (i)
    We propose a novel framework that assesses the stability of individual components within LLMs, quantifying the impact of small perturbations on the overall system.
  2. (ii)
    We conduct experiments on multiple datasets spanning vision-language models and large language models. The results consistently demonstrate the effectiveness of our approach in identifying both external and internal vulnerabilities.
  3. (iii)
    We leverage stability measure to design a parameter protection mechanism that improves the performance of various model merging techniques, demonstrating the practical value of our stability measure. algorithms.
2. Related Work
LLM Jailbreak
Prior work on LLM jailbreaking has mainly focus on the power of adversarially optimizing toward a prompt that elicits a desired LLM response (Zou et al., 2023; Fort, 2023; Liao and Sun, 2024; Geiping et al., 2024). Previous studies (Yuan et al., 2023; Huang et al., 2023) have highlighted that specific prompts, suffixes, and even so-called "magic code" (Liu et al., 2023) can induce inappropriate or harmful behaviors in language models. Recently (Fort, 2023) have found that the attack can be simplified to learning perturbation vectors added to the token embeddings, which eliminates the need to optimize over discrete tokens. (Liu et al., 2024c; Geiping et al., 2024) show similar results that manipulates embeddings could elicit refusal behavior or incorrect answers from an instruction-tuned LLM. However, the aforementioned approaches mainly focus on jailbreaking methodologies and rely on empirically specifying the target dimensions for perturbation.
VLMs Jailbreak
Despite achieving impressive performance, VLMs still face issues of adversarial robustness (Szegedy, 2013). Multiple attempts have been conducted to study the robustness of VLMs from different aspects. (Zhao et al., 2024) evaluates the adversarial robustness of MLLMs on image captioning under white-box settings, while (Schlarmann and Hein, 2023) conducts both transfer-based and query-based attacks on MLLMs assuming black-box access. (Carlini et al., 2024; Qi et al., 2023) trigger VLMs to generate toxic content by imposing adversarial perturbations to the input images. (Bailey et al., 2023) studies image hijacks to achieve specific string, leak context, and jailbreak attacks. These exploratory works demonstrate that MLLMs still face stability and security issues under adversarial perturbations.
3. Local Stability Measure of Large language models
@@
In this section, we propose a new metric called FI to quantify the stability of large language models against local perturbations. Specifically, we focus on auto-regressive language models, which generate one token at a time based on both the initial input and previously generated tokens. Considering this sequential generation nature, we first develop FI for single-step token generation tasks and discuss its theoretical and computational properties in great detail. Following this, we demonstrate how the proposed stability measure can be naturally extended to address sequence generation tasks.
@@
FI metric.
Since our primary interest lies in examining the behavior of $P(y|x,\omega,\theta)$ as a function of $\omega$ near $\omega_{0}=0$, we shift focus from $\theta$ to $\omega$. We introduce the perturbation manifold as defined in (Zhu et al., 2007) and (Zhu et al., 2011).
@@
The metric $g_{\omega}$ on $\mathcal{M}$ can be defined with the metric tensor $G_{\omega}$. Consider two tangent vectors at $\omega$ given by
$v_{j}(\omega)=h_{j}^{\top}\partial_{\omega}\ell(\omega|y,x,\theta)\in T_{\omega},\quad j=1,2$
where $h_{1}$ $h_{2}$ are the weights on the basis. Their inner product is defined as:
$\langle v_{1}(\omega),v_{2}(\omega)\rangle_{g_{\omega}}=\sum_{y=1}^{K}v_{1}(\omega)v_{2}(\omega)P(y|x,\theta)=h_{1}^{\top}G_{\omega}h_{2}.$
@@
Subsequently, the norm of $v_{j}(\omega)$ under metric $g_{\omega}$ is:
$\|v_{j}\|_{g_{\omega}}=\sqrt{h_{j}^{\top}G_{\omega}h_{j}}.$
Let $C(t)=P(y|x,\theta,\omega(t))$ be a smooth curve on the manifold $\mathcal{M}$ connecting two points $\omega_{1}=\omega(t_{1})$ and $\omega_{2}=\omega(t_{2})$. Then, the distance between $\omega_{1}$ and $\omega_{2}$ along the curve $C(t)$ is given by:
@@
(2) $\mathbf{FI}_{{\omega}}\left(\omega_{0}\right)=\max_{C}\lim_{t\rightarrow 0}\frac{[f({\omega}(t))-f({\omega}(0))]^{2}}{S_{C}^{2}({\omega}(t),{\omega}(0))}.$

### v2

@@
Abstract
Large Language Models (LLMs) and Vision-Language Models (VLMs) have achieved impressive performance across a wide range of tasks, yet they remain vulnerable to carefully crafted perturbations. In this study, we seek to pinpoint the sources of this fragility by identifying parameters and input dimensions (pixels or token embeddings) that are susceptible to such perturbations. To this end, we propose a stability measure called FI, First order local Influence, which is rooted in information geometry and quantifies the sensitivity of individual parameter and input dimensions. Our extensive analysis across LLMs and VLMs (from 1.5B to 13B parameters) reveals that: (I) A small subset of parameters or input dimensions with high FI values disproportionately contribute to model brittleness. (II) Mitigating the influence of these vulnerable parameters during model merging leads to improved performance.
Breach in the Shield: Unveiling the Vulnerabilities of Large Language Models
Runpeng Dai1∗ Run Yang2∗  Fan Zhou3  Hongtu Zhu1† 1University of North Carolina at Chapel Hill  2BiliBili 3Shanghai University of Finance and Economics {runpeng, htzhu}@email.unc.edu yangrun@bilibili.com zhoufan@mail.shufe.edu.cn
1 Introduction
 [^1] Both authors contributed equally to this work.
Large Language Models (LLMs) and Vision Language Models (VLMs) such as GPT (Brown et al., 2020) and Llama (Touvron et al., 2023), have revolutionized the field of Natural Language Processing (NLP), exhibiting remarkable proficiency across a variety of tasks (Gong et al., 2024; Zheng et al., 2025b; Luo et al., 2025) and modalities (Bai et al., 2023; Liu et al., 2024b; Zheng et al., 2025a). These modern LLMs are massive in size, trained on vast amounts of data, and meticulously aligned to prevent from generating harmful content (Perez et al., 2022), leaking private information (Zhang et al., 2024), or exhibiting sexual or religious bias (Xie and Lukasiewicz, 2023).
@@ 1 Introduction
Externally, LLMs are vulnerable to input perturbations, such as Embedding-Corrupted Prompts (Fort, 2023; Liu et al., 2024a). This susceptibility extends to visual inputs in VLMs, where adversarially optimized images can drastically alter model behavior (Qi et al., 2024). Beyond adversarial attacks, VLMs exhibit high sensitivity to perturbations in specific local regions of an image—a common issue, as user-uploaded images often suffer from blurring, masking, or low resolution. The vulnerability is highlighted in our case study of the Qwen-VL model. As depicted in Figure 1, masking the ten most sensitive pixels, which are unrelated to the question, resulted in incorrect model outputs.
Figure 1: A case study of the Qwen-VL model (Bai et al., 2023) on SCI-QA. The image on the far right visualizes the per-pixel FI values. Masking just 10 pixels with the highest FI values leads to a failure in producing the correct answer.
@@
Internally, LLM stability is further challenged by parameter perturbations, often introduced through model merging and quantization. While these techniques improve deployment efficiency by reducing inference costs (Frantar and Alistarh, 2023; Ashkboos et al., 2024), they can also induce hallucinations and degrade performance (Men et al., 2024; Yu et al., 2024; Li et al., 2024). However, our findings reveal that parameter susceptibility varies significantly. As Figure 2 illustrates, randomly dropping 5% of parameters has a minimal impact on performance. In contrast, zeroing out just 1% of the parameters identified by our measure can drastically reduce accuracy, even below random guessing levels.
Figure 2: A case study of Qwen2 on MMLU-geography. This figure illustrates the decline in response accuracy as a greater percentage of parameters are zeroed. “FI-High” targets parameters with the highest FI values, while “Random” indicates random parameter removal.
@@
To pinpoint the sources of this fragility, we propose a novel stability measure called FI, First order local Influence, to quantitatively assess the stability of LLMs against perturbations. Specifically, we construct a perturbation manifold that encompasses all perturbed models, along with its associated geometric properties. Our stability measure quantifies the degree of local influence of a perturbation on a given objective function within this manifold, thereby reflecting the stability of individual LLM components.
@@
FI’s versatility allows for effective stability assessment under both external and internal perturbations across various granularities—from individual parameters to input features like pixels and patches.
@@
FI effectively identifies vulnerabilities. Our extensive studies validate its effectiveness in pinpointing fragile pixels in VLM vision inputs, vulnerable embedding dimensions of tokens in LLMs (Section 4), and salient model parameters (Subsection 5.1).
FI offers insights into improving model robustness. We further illustrate that understanding these vulnerabilities can lead to enhanced model resistance to perturbations. By focusing on model merging as an example, we show that safeguarding key parameters identified by high FI values can substantially reduce performance degradation during the merging process (Subsection 5.2).
2 Related Work
Recent efforts to evaluate LLM stability typically adopt a coarse-grained approach, aiming to assess the overall robustness of models under various perturbations. One line of work investigates how stability is influenced by sampling parameters, such as temperature, which affect output variability during generation (Atil et al., 2024; Ouyang et al., 2025). Another direction studies model sensitivity to input or parameter perturbations. For instance, Liu et al. analyze input-level robustness using optimal transport to quantify a model’s response to distributional shifts in prompts. On the other hand, Peng et al. (2024) focus on parameter-space perturbations, demonstrating that LLMs remain robust to weight changes up to a certain threshold, beyond which performance significantly degrades. They estimate a model’s robustness tolerance by injecting random perturbations into model weights and evaluating the performance drop.
Despite these contributions, fine-grained analyses—such as those examining the effect of individual input tokens, pixels, or specific model parameters—remain underexplored. Wei et al. (2024) take a step in this direction by leveraging pruning-based techniques, including SNIP (Lee et al., 2018) and Wanda (Sun et al., 2023), to identify critical neurons and low-rank structures that impact model safety and utility. However, there is still a lack of unified metrics or frameworks that assess stability with respect to both input- and parameter-level perturbations. Moreover, the downstream applications of such stability assessments remain largely unexamined.
3 Stability Measure of Large language models
@@
In this section, we propose a new metric called FI to quantify the stability of large language models against local perturbations. Considering the auto-regressive nature of LLMs, we first develop FI for single-step generation and discuss its theoretical and computational properties in detail. We then show how FI can be naturally extended to sequence generation tasks. Finally, we compare FI to existing stability measures, highlighting its unique advantages.
3.1 FI Metric
@@
Perturbation Manifold and FI
Since our primary interest lies in examining the behavior of $P(y|x,\omega,\theta)$ as a function of $\omega$ near $\omega_{0}=0$, we shift focus from $\theta$ to $\omega$. We introduce the perturbation manifold as defined in Zhu et al. (2007) and Zhu et al. (2011).
@@
The metric $g_{\omega}$ on $\mathcal{M}$ can be defined with the metric tensor $G_{\omega}$. Consider two tangent vectors at $\omega$ given by $v_{j}(\omega)=h_{j}^{\top}\partial_{\omega}\ell(\omega|y,x,\theta)\in T_{\omega}$, where $h_{1}$ $h_{2}$ are the weights on the basis. Their inner product is defined as:
$\langle v_{1}(\omega),v_{2}(\omega)\rangle_{g_{\omega}}=\sum_{y=1}^{K}v_{1}(\omega)v_{2}(\omega)P(y|x,\theta).$
@@
Subsequently, the norm of $v_{j}(\omega)$ under metric $g_{\omega}$ is $\|v_{j}\|_{g_{\omega}}=\sqrt{h_{j}^{\top}G_{\omega}h_{j}}$. Let $C(t)=P(y|x,\theta,\omega(t))$ be a smooth curve on the manifold $\mathcal{M}$ connecting two points $\omega_{1}=\omega(t_{1})$ and $\omega_{2}=\omega(t_{2})$. Then, the distance between $\omega_{1}$ and $\omega_{2}$ along the curve $C(t)$ is given by:
@@
$\mathbf{FI}_{{\omega}}\left(\omega_{0}\right)=\max_{C}\lim_{t\rightarrow 0}\frac{[f({\omega}(t))-f({\omega}(0))]^{2}}{S_{C}^{2}({\omega}(t),{\omega}(0))}.$ (1)